# Determination of the stably free cancellation property for orders

### v1

@@ 12.3. Computation of $S$
is exact, it suffices to find generators for the first and third group. Generators for $\frac{1+\mathfrak{Q}_{1}}{1+\mathfrak{Q}}$ can be computed as in [BB06, §3.7]. For the computation of generators of the non-trivial group on the right hand side of (12.4), the elements can be listed as follows: determine the set of representatives of $\tilde{\mathfrak{Q}}/\mathfrak{Q}_{1}$ and check for each such representative $\alpha$ whether $1+\alpha$ it is a unit mod $\mathfrak{Q}$.
@@ 12.4. A probabilistic algorithm
Suppose that $A_{1}$ is Eichler and that every simple component of $A_{2}$ is a totally definite quaternion algebra. Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. If the following algorithm returns $\mathsf{true}$, then $X$ is free over $\Lambda$; if returns $\mathsf{inconclusive}$, then nothing can be said about whether $X$ is free over $\Lambda$ or not.
@@
Apart from Step (2), all steps are the essentially the same as in Algorithm 12.12. Step (2) can be performed as follows. Recall that $S=\overline{\mathrm{nr}}(\overline{\pi}_{1}(\Sigma))$. Thus in order to find a subgroup $S^{\prime}$ of $S$ it is sufficient to find a subgroup of
@@
Random elements of $\Sigma$ can be found by sampling preimages of $\overline{1}\in\Lambda_{2}/\mathfrak{f}_{2}$ under the canonical map $\Lambda/\mathfrak{f}\longrightarrow\Lambda_{2}/\mathfrak{f}_{2}$ (a homomorphism of finitely generated abelian groups) and discarding the non-invertible elements. In this way we obtain a sequence of elements $(\overline{\alpha}_{i})_{i\in\mathbb{Z}_{\geq 1}}$ in $\Sigma$ and determine $S^{\prime}:=\langle\overline{\mathrm{nr}}(\bar{\pi}_{1}(\bar{\alpha}_{i}))\mid i\leq B\rangle$, where $B\in\mathbb{Z}_{\geq 1}$ is either a fixed predetermined bound or chosen dynamically until $S^{\prime}$ stabilises. ∎
@@ 13. Further results on SFC for integral group rings
$\tilde{T}\times C_{2},\,\tilde{T}\times Q_{12},\,\tilde{T}\times Q_{20},\,Q_{8}\rtimes\tilde{T},\,G_{(96,66)},\,G_{(192,183)},\,G_{(480,962)}.$
@@
Proof.
@@
For $G=G_{(96,66)}$, $\tilde{T}\times Q_{12}$ or $\tilde{T}\times Q_{20}$, there exists a normal subgroup $N$ of order $2$, such that $H:=G/N$ is isomorphic to $\mathrm{GL}_{2}(\mathbb{F}_{3})$, $G_{(144,127)}$, or $G_{(280,108)}$, respectively. In each case, $\mathbb{Q}[H]$ satisfies the Eichler condition and hence so does $\mathbb{Q}[H\times C_{2}]$. Thus $\mathbb{Z}[H\times C_{2}]$ has LFC by Theorem 3.4. Let $\Gamma=\mathbb{Z}[G]/\mathrm{Tr}_{N}\mathbb{Z}[G]$. The cardinalities of the locally free class groups of $\mathbb{Z}[H]$, $\mathbb{Z}[G]$, $\mathbb{Z}[H\times C_{2}]$ and $\Gamma$ can be computed using the the algorithm of [BB06, §3], and thus it can be verified that Proposition 6.10 can be applied to show that $\mathbb{Z}[G]$ has SFC if only if $\Gamma$ has SFC.
@@
For $G=Q_{8}\rtimes\tilde{T}\cong G_{(192,1022)}$, there is a normal subgroup $N$ of order $2$ such that $H:=G/N$ is isomorphic to $G_{(96,204)}$. It can easily be checked that $\mathbb{Q}[H]$ satisfies the Eichler condition. Let $\Gamma=\mathbb{Z}[G]/\mathrm{Tr}_{N}\mathbb{Z}[G]$. Unfortunately, Proposition 6.10 cannot be applied in this situation because the required equality (6.3) does not hold. However, $\mathbb{Z}[G]$ can be shown to have SFC if and only if $\Gamma$ has SFC by an application of Algorithm 10.3 with $\Lambda_{1}=\mathbb{Z}[H]$ and $\Lambda_{2}=\Gamma$. Note that $\overline{\Lambda}=\mathbb{F}_{2}[H]$, and so an explicit embedding of $\overline{\Lambda}^{\times}$ into $\mathrm{GL}_{96}(\mathbb{F}_{2})$ can be determined as explained in Remark 6.3 (see also §11.3). Crucially, this allows computations in $\overline{\Lambda}^{\times}$ to be performed reasonably efficiently.
@@
Therefore, for $G=G_{(96,66)}$, $G_{(192,1022)}$, $\tilde{T}\times Q_{12}$, or $\tilde{T}\times Q_{20}$, we have shown that $\mathbb{Z}[G]$ has SFC if and only if $\Gamma$ has SFC. This is highly advantageous from a computational point of view, since $\dim_{\mathbb{Q}}\mathbb{Q}\otimes_{\mathbb{Z}}\Gamma=\frac{1}{2}\dim_{\mathbb{Q}}\mathbb{Q}[G]=\frac{1}{2}|G|$. To verify that $\Gamma$ has SFC, we again apply Algorithm 10.3, this time with the fibre product induced by the decomposition $\mathbb{Q}\otimes_{\mathbb{Z}}\Gamma=B_{1}\oplus B_{2}$, with $B_{1}$ the largest component of $\mathbb{Q}\otimes_{\mathbb{Z}}\Gamma$ satisfying the Eichler condition (equivalently, with $B_{2}$ the largest component that is the direct sum of totally definite quaternion algebras). Computations in $\overline{\Lambda}^{\times}$ are performed by first computing an explicit embedding into the symmetric group $\mathfrak{S}_{d}$ for some $d\in\mathbb{Z}_{>0}$ (see §11.3). This is less efficient than working with an embedding into $\mathrm{GL}_{n}(\mathbb{F}_{p})$ for appropriate choices of $n$ and $p$, but such an embedding does not exist in general.
@@
For $G=G_{(192,183)}$ or $G_{(480,962)}$, there exists a normal subgroup $N$ of such that $H:=G/N\cong G_{(96,66)}$ and so $\mathbb{Z}[H]$ has SFC by the result for $G_{(96,66)}$. Let $\Gamma=\mathbb{Z}[G]/\mathrm{Tr}_{N}\mathbb{Z}[G]$. It can easily be checked that $\mathbb{Q}\otimes_{\mathbb{Z}}\Gamma$ satisfies the Eichler condition and thus $\Gamma$ has LFC by Theorem 3.4. Then $\mathbb{Z}[G]$ can be shown to have SFC by an application of Algorithm 10.3 with $\Lambda_{1}=\Gamma$ and $\Lambda_{2}=\mathbb{Z}[H]$. Note that $\overline{\Lambda}=\mathbb{F}_{p}[H]$ for $p=2$ or $5$, which allows computations in $\overline{\Lambda}^{\times}$ to be performed reasonably efficiently, as explained above. ∎
@@
    $\tilde{T}\times C_{2}^{2}$, $\tilde{O}\times C_{2}$, $\tilde{I}\times C_{2}$,
@@
We use the GAP [GAP24] small groups database [BEO02], which in particular, contains all finite groups $G$ with $|G|\leq 1023$. For each such $G$, we first check whether $\mathbb{Z}[G]$ has SFC by Theorems 4.5, 4.10, 4.11, and 13.1, as well as Corollary 6.9. We then check whether $G$ has a quotient isomorphic to any of the groups listed in (i), (ii), (iii), or (iv), in which case $\mathbb{Z}[G]$ fails SFC by Corollary 9.7. These checks can easily be implemented as a program in GAP [GAP24] or Magma [BCP97]. For $G_{(480,256)}\cong C_{5}\times G_{(96,66)}$, $G_{(576,7442)}\cong C_{3}\times(Q_{8}\rtimes\tilde{T})$, and $G_{(960,9669)}\cong C_{5}\times(Q_{8}\rtimes\tilde{T})$, combination of Theorems 4.6 and 13.1 shows that $\mathbb{Z}[G]$ has SFC. The remaining groups consist of all those listed in (a)–(e). ∎
@@
    $\tilde{T}\times C_{2}^{2}$, $\tilde{O}\times C_{2}$, $\tilde{I}\times C_{2}$,

### v2

@@ 12.3. Computation of $S$
is exact, it suffices to find generators for the first and third group. Generators for $\frac{1+\mathfrak{Q}_{1}}{1+\mathfrak{Q}}$ can be computed as in [BB06, §3.7]. For the computation of generators of the non-trivial group on the right hand side of (12.4), the elements can be listed as follows: determine a set of representatives of $\tilde{\mathfrak{Q}}/\mathfrak{Q}_{1}$ and check for each such representative $\alpha$ whether $1+\alpha$ it is a unit mod $\mathfrak{Q}$.
@@ 12.4. A probabilistic algorithm
Suppose that $A_{1}$ satisfies the Eichler condition and that every simple component of $A_{2}$ is a totally definite quaternion algebra. Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. If the following algorithm returns $\mathsf{true}$, then $X$ is free over $\Lambda$; if it returns $\mathsf{inconclusive}$, then nothing can be said about whether $X$ is free over $\Lambda$ or not.
@@
Apart from Step (2), all steps are essentially the same as in Algorithm 12.5. Step (2) can be performed as follows. Recall that $S=\overline{\mathrm{nr}}(\overline{\pi}_{1}(\Sigma))$. Thus in order to find a subgroup $S^{\prime}$ of $S$ it is sufficient to find a subgroup of
@@
Random elements of $\Sigma$ can be found by sampling preimages of $\overline{1}\in\Lambda_{2}/\mathfrak{f}_{2}$ under the canonical map $\Lambda/\mathfrak{f}\longrightarrow\Lambda_{2}/\mathfrak{f}_{2}$ (a homomorphism of finitely generated abelian groups) and discarding the non-invertible elements. In this way we obtain a sequence of elements $(\overline{\alpha}_{i})_{i\in\mathbb{Z}_{\geq 1}}$ in $\Sigma$ and determine $S^{\prime}:=\langle\overline{\mathrm{nr}}(\overline{\pi}_{1}(\overline{\alpha}_{i}))\mid i\leq B\rangle$, where $B\in\mathbb{Z}_{\geq 1}$ is either a fixed predetermined bound or chosen dynamically until $S^{\prime}$ stabilises. ∎
@@ 13. Further results on SFC for integral group rings
$\tilde{T}\times C_{2},\,\tilde{T}\times Q_{12},\,\tilde{T}\times Q_{20},\,Q_{8}\rtimes\tilde{T},\,Q_{8}\rtimes Q_{12},\,\tilde{I}\times C_{2},\,G_{(192,183)},\,G_{(384,580)},\,G_{(480,962)}.$ (13.1)
@@
Remark 13.2.
In order, the first six groups listed in (13.1) are isomorphic to
$G_{(48,32)},\,G_{(288,409)},\,G_{(480,266)},\,G_{(192,1022)},\,G_{(96,66)},\,G_{(240,94)}.$
Proof of Theorem 13.1.
@@
Let $G$ be any group listed in (13.1) and let $\mathcal{M}$ be any maximal $\mathbb{Z}$-order in $\mathbb{Q}[G]$. Then every binary polyhedral quotient of $G$ is isomorphic to $Q_{12}$, $Q_{20}$, $\tilde{T}$, $\tilde{O}$, or $\tilde{I}$, and thus $\mathcal{M}$ has LFC by Corollary 4.6. Hence every direct summand of $\mathcal{M}$ has LFC by Corollary 3.3. Thus in all applications of Algorithm 10.3 below, it is not necessary to perform Step (3) as the maximal orders in question are already known to have LFC (and thus SFC).
@@
For $G=Q_{8}\rtimes Q_{12}$, $\tilde{T}\times Q_{12}$, $\tilde{T}\times Q_{20}$, or $\tilde{I}\times C_{2}$, there exists a normal subgroup $N$ of order $2$, such that $H:=G/N$ is isomorphic to $\mathrm{GL}_{2}(\mathbb{F}_{3})$, $G_{(144,127)}$, $G_{(240,108)}$, or $C_{2}\times A_{5}$, respectively. In each case, $\mathbb{Q}[H]$ satisfies the Eichler condition and this implies that $\mathbb{Q}[H\times C_{2}]$ does as well. Thus $\mathbb{Z}[H\times C_{2}]$ has LFC by Theorem 3.4. Let $\Gamma=\mathbb{Z}[G]/\mathrm{Tr}_{N}\mathbb{Z}[G]$. The cardinalities of the locally free class groups of $\mathbb{Z}[H]$, $\mathbb{Z}[G]$, $\mathbb{Z}[H\times C_{2}]$ and $\Gamma$ can be computed using the the algorithm of [BB06, §3], and thus it is easily verified that equality (6.3) holds. Therefore Proposition 6.10 shows that $\mathbb{Z}[G]$ has SFC if only if $\Gamma$ has SFC.
@@
For $G=Q_{8}\rtimes\tilde{T}$, there is a normal subgroup $N$ of order $2$ such that $H:=G/N$ is isomorphic to $G_{(96,204)}$. It can easily be checked that $\mathbb{Q}[H]$ satisfies the Eichler condition. Let $\Gamma=\mathbb{Z}[G]/\mathrm{Tr}_{N}\mathbb{Z}[G]$. Unfortunately, Proposition 6.10 cannot be applied in this situation because the required equality (6.3) does not hold. However, $\mathbb{Z}[G]$ can be shown to have SFC if and only if $\Gamma$ has SFC by an application of Algorithm 10.3 with $\Lambda_{1}=\mathbb{Z}[H]$ and $\Lambda_{2}=\Gamma$. Note that $\overline{\Lambda}=\mathbb{F}_{2}[H]$, and so an explicit embedding of $\overline{\Lambda}^{\times}$ into $\mathrm{GL}_{96}(\mathbb{F}_{2})$ can be determined as explained in Remark 6.3 (see also §11.3). Crucially, this allows computations in $\overline{\Lambda}^{\times}$ to be performed reasonably efficiently.
@@
Therefore, for $G=Q_{8}\rtimes Q_{12}$, $Q_{8}\rtimes\tilde{T}$, $\tilde{T}\times Q_{12}$, $\tilde{T}\times Q_{20}$, or $\tilde{I}\times C_{2}$, we have shown that $\mathbb{Z}[G]$ has SFC if and only if $\Gamma$ has SFC. This is highly advantageous from a computational point of view, since $\dim_{\mathbb{Q}}\mathbb{Q}\otimes_{\mathbb{Z}}\Gamma=\frac{1}{2}\dim_{\mathbb{Q}}\mathbb{Q}[G]=\frac{1}{2}|G|$. To verify that $\Gamma$ has SFC, we again apply Algorithm 10.3, this time with the fibre product induced by the decomposition $\mathbb{Q}\otimes_{\mathbb{Z}}\Gamma=B_{1}\oplus B_{2}$, with $B_{1}$ the largest component of $\mathbb{Q}\otimes_{\mathbb{Z}}\Gamma$ satisfying the Eichler condition (equivalently, with $B_{2}$ the largest component that is the direct sum of totally definite quaternion algebras). Computations in $\overline{\Lambda}^{\times}$ are performed by first computing an explicit embedding into the symmetric group $\mathfrak{S}_{d}$ for some $d\in\mathbb{Z}_{>0}$ (see §11.3). This is less efficient than working with an embedding into $\mathrm{GL}_{n}(\mathbb{F}_{p})$ for appropriate choices of $n$ and $p$, but such an embedding does not exist in general.
@@
For $G=G_{(192,183)}$ or $G_{(480,962)}$, there exists a normal subgroup $N$ of $G$ such that $H:=G/N$ is isomorphic to $Q_{8}\rtimes Q_{12}$ and so $\mathbb{Z}[H]$ has SFC by the result for $Q_{8}\rtimes Q_{12}$. Let $\Gamma=\mathbb{Z}[G]/\mathrm{Tr}_{N}\mathbb{Z}[G]$. It can easily be checked that $\mathbb{Q}\otimes_{\mathbb{Z}}\Gamma$ satisfies the Eichler condition and thus $\Gamma$ has LFC by Theorem 3.4. Then $\mathbb{Z}[G]$ can be shown to have SFC by an application of Algorithm 10.3 with $\Lambda_{1}=\Gamma$ and $\Lambda_{2}=\mathbb{Z}[H]$. Note that $\overline{\Lambda}=\mathbb{F}_{p}[H]$ for $p=2$ or $5$, which allows computations in $\overline{\Lambda}^{\times}$ to be performed reasonably efficiently, as explained above.
For $G=G_{(384,580)}$ there exists a normal subgroup $N$ of $G$ such that $H:=G/N$ is isomorphic to $G_{(192,183)}$ and so $\mathbb{Z}[H]$ has SFC by the result for $G_{(192,183)}$. Then the same method as in the previous paragraph with $\overline{\Lambda}=\mathbb{F}_{2}[H]$ shows that $\mathbb{Z}[G]$ has SFC. ∎
Remark 13.3.
The computations involved in the proof of Theorem 13.1 rely on the algorithm of [BB06, §3] to compute locally free class groups and on the new freeness testing methods of §12. These in turn rely on the computation of ray class groups and of unit groups of rings of integers of number fields. Due to reasons of efficiency, state-of-the-art algorithms to perform these computations yield solutions that are provably correct under the assumption that the Generalized Riemann Hypothesis (GRH) holds. In the proof of Theorem 13.1, the number fields involved are the character fields of the irreducible complex characters of the finite groups $G$ listed in (13.1):
$\mathbb{Q},\,\mathbb{Q}(i),\,\mathbb{Q}(\sqrt{-3}),\,\mathbb{Q}(\sqrt{5}),\,\mathbb{Q}(\sqrt{2}),\,\mathbb{Q}(\sqrt{-2}),\,\mathbb{Q}(\zeta_{8}),\,\mathbb{Q}(\sqrt{-15}),\,\mathbb{Q}(i,\sqrt{3}),\,\mathbb{Q}(\sqrt{-3},\sqrt{5}).$
As these fields are of small degree and discriminant, the necessary computations can be performed unconditionally (or can be found in the literature) and hence Theorem 13.1 does not depend on GRH.
@@
    $\tilde{T}\times C_{2}^{2}$, $\tilde{O}\times C_{2}$, $\tilde{I}\times C_{2}^{2}$,
@@
We use the GAP [GAP24] small groups database [BEO02], which in particular, contains all groups $G$ with $|G|\leq 1023$. For each such $G$, we perform the following checks, which can easily be implemented as a program in GAP [GAP24] or Magma [BCP97].
  • If $G$ has a quotient isomorphic to any of the groups listed in (i), (ii), (iii), or (iv), then $\mathbb{Z}[G]$ fails SFC by Corollary 9.8.
  • If $m_{\mathbb{H}}(G)\leq 1$ then $\mathbb{Z}[G]$ has SFC by Theorem 4.13.
  • If $G$ has a binary polyhedral quotient $H$ listed in Theorem 4.2 such that $m_{\mathbb{H}}(G)=m_{\mathbb{H}}(H)$ then $\mathbb{Z}[G]$ has SFC by Theorem 4.12.
  • If $G$ has a quotient isomorphic to $\tilde{T}\times C_{2}$ and $m_{\mathbb{H}}(G)=2$ then $\mathbb{Z}[G]$ has SFC by Corollary 6.9.
  • If $G\cong G_{(576,5128)}\cong\tilde{T}\times\tilde{T}$ then $\mathbb{Z}[G]$ has SFC by Theorem 4.7.
  • If $G\cong H\times K$ where $H$ is any group listed in Theorem 13.1 and $K$ is any of $C_{3},C_{5},C_{7},C_{9}$ or $C_{3}^{2}$, then $\mathbb{Z}[G]$ has SFC by Theorems 13.1 and 4.8.
The groups $G$ with $|G|\leq 1023$ that do not satisfy any of the above conditions are listed in (a)–(g) above. ∎
The following result is Theorem 1.4 from the introduction.
@@
    $\tilde{T}\times C_{2}^{2}$, $\tilde{O}\times C_{2}$, $\tilde{I}\times C_{2}^{2}$,